OpenStudy (anonymous):

What is the concentration in mo/dm^3, of a solution containing:4.25g silver nitrate,AgNO3, in 500cm^3 solution

OpenStudy (anonymous):

concentration=moles/litres mol.wt.of AgNO3=108+14+3*16=170 moles of Agno3=4.25/170=0.025 10^6 cm3=10^3L 500 cm3=0.5L concentration=0.025/0.5=0.05mol/lit is it right???

OpenStudy (anonymous):

Sarkar's answer is all correct but you asked for it in moles per deciliter. There are tenten deciliters for every one liter. Deci means one tenth. So if you have .5L then you have 5 deciliters. Therefore the concentration is \[.5L \times {10dL \over 1L} \]\[{.025moles \over 5dL} = .005M\]

OpenStudy (anonymous):

ahhh,thanks @shamalamadingdong

OpenStudy (anonymous):

yeah thanks folks :)

OpenStudy (anonymous):

No problem.

OpenStudy (mos1635):

???? 1 Lt = ! dm^3 isn't that so????

OpenStudy (anonymous):

yes @mos1635